The shape of the distribution of the time required to get an oil change at a 15​-minute ​oil-change facility is unknown.​ However, records indicate that the mean time is 16.4 minutes​, and the standard deviation is 3.8 minutes. Complete parts ​(a) through ​(c).

(c) Suppose the manager agrees to pay each employee a​ $50 bonus if they meet a certain goal. On a typical​ Saturday, the​ oil-change facility will perform 45 oil changes between 10 A.M. and 12 P.M. Treating this as a random​ sample, there would be a​ 10% chance of the mean​ oil-change time being at or below what​ value? This will be the goal established by the manager. There is a​ 10% chance of being at or below a mean​ oil-change time of nothing minutes. ​(Round to one decimal place as​ needed.)
Answer #1

Solution: Given that mean = 16.4, sd = 3.8, n = 45

(c) P(X < x) = 0.1 for Z = -1.2816
  
X = mean +Z*sd/sqrt(n) = 16.4 - 1.2816*3.8/sqrt(45)
= 15.674
= 15.7 (rounded)